# ScaffoldGPT: A Scaffold-based Large Language Model for Drug Improvement

### v1

@@ 4 The ScaffoldGPT Algorithm
[…]
Stage 1. Pretrain a LLM-based generator. Let us note the LLM-based generator policy as $\pi_{\theta}$, which computes the probability $p$ of the occurrence of the $t^{th}$ token in a target molecule $Y$. It takes into account all preceding tokens $\mathbf{y}_{<t}=\left[y_{1},...,y_{t-1}\right]$ in the target, as well as the scaffold compound $S$, which is noted as $\pi_{\theta}\left({y_{t}{\,|\,}\mathbf{y}_{<t},S}\right)=p\left({y_{t}|\mathbf{y}_{<t},S}\right).$ The parameters $\theta$ of the generator policy $\pi_{\theta}$ are trained using the training corpus set through the minimization of the negative log-likelihood (NLL) for the complete SMILES strings across the entire set. This process is described as follows:
$NLL$ $=-\log{P}\left({Y|S}\right)=-\sum_{t=1}^{T}\log{P}\left({y_{t}|y_{t-1},...,y_{1},S}\right)$
$=-\sum_{t=1}^{T}\log\pi_{\theta}\left({y_{t}|y_{1:t-1},S}\right),$ (3)
In this project, we employ pre-training to harness large quantities of unlabeled text to construct a basic foundation model of language understanding. This foundation model can subsequently be fine-tuned and tailored to meet various specialized goals. In this work, we propose a novel framework for pre-training a Transformer and linking scaffolds with complete molecules, based on a SMILES (Simplified Molecular Input Line Entry System) [47] string representation of the molecule. Here, we define $\mathcal{S}$ as the scaffold space and $\mathcal{Y}$ as the target molecule space. With $\mathcal{P}=\left\{\left({s,y}\right)|s,y\in\mathcal{S}\times\mathcal{Y}\right\}$, we denote the set of scaffold and molecular pairs from $\mathcal{S}$ and $\mathcal{Y}$. In this notation, $s$ represents the scaffold of the target molecule, and $y$ is the corresponding target molecule. We initially pre-trained our tokenizer using the Byte Pair Encoding (BPE) method [11, 38]. Building on the pre-trained BPE tokenizer, we propose a two-phase incremental training approach, as illustrated in Fig. 1, to notably enhance the model’s ability to improve the validity of inferring the target molecule from its scaffold.
Incremental training.
The rationale for incremental training is to conduct local optimization before embarking on global optimization. Therefore, we divide our training into two phases. In the first phase, we focus on training a large language model (LLM) exclusively for molecules using Causal Language Modeling (CLM). CLM utilizes an autoregressive method where the model is trained to predict the next token in a sequence by considering only the tokens that precede it. The phase 1 corpus is designed as follows:
$C_{\text{Phase 1}}=\left\{[BOS],\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},$ (4)
In the second phase, building upon the success of the LLM model developed in phase 1, which demonstrated high accuracy in molecular generation, we advance the training by focusing on pairs of scaffolds and molecules using CLM. The phase 2 corpus is designed as follows:
$C_{\text{Phase 2}}=\left\{\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}}\right\},$ (5)
Consequently, the model can generate the appropriate molecule when given a scaffold. However, since a single scaffold may correspond to multiple molecules, we further refine the LLM-based generator policy, $\pi_{\theta}$, to target specific outcomes by applying reinforcement learning finetuning in the next stage.
Stage 2. RL finetuning. Fine-tuning a generative model is crucial for producing outcomes that meet specific objectives. In this study, we use the Advantage-alignment Policy Optimization (APO) [22] algorithm to fine-tune the pretrained LLM-based generator policy $\pi_{\theta}$. This approach steers the model from a given scaffold towards the targeted molecule, simultaneously improving multiple properties.
In this work, we adopt the define of reward function from Liu et al. [23], and regarded each pharmaceutical property as a critic $C$ and got an ensemble critics $\mathbf{C}$ as follows:
${R^{\text{norm}}\left({Y}\right)}:={R^{\text{norm}}\left({Y|S}\right)}$ (6)
$=\lambda\cdot\text{Norm}\left({{C}^{\text{Tanimoto}}\left({S,Y}\right)}\right)+\sum_{i=0}^{|\mathbf{C}|-1}\lambda\cdot\text{Norm}\left({{C}_{i}{\left({Y}\right)}}\right),$
where Norm is employed to standardize diverse attributes to a consistent scale. In this instance, Norm refers to the process of min-max normalization, which is used to adjust the attributes so they fit within the $\left[0,1\right]$ range.
In this work, we use BON [12] (Best of N) search to estimate the reward for a prompt (partial molecule). BON can be formulated as the following:
where $\text{where }Y_{j}=\left[\mathbf{y}_{<i},y_{i},\cdots,y_{T}\right]_{j},$
$\text{and }y_{i}\sim\textsc{Top-PK}\left({\mathbf{y}_{<i},p,k}\right)|_{S}.$
where Top-PK [23] defined as follows:
[…]
 $\text{~{}~{}~{}~{}~{}~{}~{}}j=\min\left\{\mathop{\rm arg\,min}_{j^{\prime}}\sum_{i=1}^{j^{\prime}}\pi_{\theta}\left({y_{i}|S,\mathbf{y}_{<i}}\right)\geq p,k\right\},$
 $\text{~{}~{}~{}~{}~{}~{}~{}}\text{and }\pi_{\theta}\left({y_{g}|S,\mathbf{y}_{<i}}\right)>\pi_{\theta}\left({y_{h}|S,\mathbf{y}_{<i}}\right),\text{if }g<h,$
@@
(PDBID: MMP [26] 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ $0$8.259 $\pm$ 0.004 0.691 $\pm$ 0.001 2.682 $\pm$ 0.004 3.109 $\pm$ 0.020 0.862 $\pm$ 0.000
7BQY) Similarity ($\geq$ 0.5) [26] 0.995 $\pm$ 0.001 0.615 $\pm$ 0.000 0.706 $\pm$ 0.001 -8.165 $\pm$ 0.024 0.697 $\pm$ 0.004 2.621 $\pm$ 0.006 3.180 $\pm$ 0.029 0.782 $\pm$ 0.001
Similarity ([0.5, 0.7)]) [26] 0.995 $\pm$ 0.001 0.612 $\pm$ 0.001 0.701 $\pm$ 0.001 -8.187 $\pm$ 0.010 0.691 $\pm$ 0.001 2.611 $\pm$ 0.009 3.240 $\pm$ 0.014 0.756 $\pm$ 0.003
Similarity ($\geq$ 0.7) [26] 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ 0.001 -8.214 $\pm$ 0.002 0.691 $\pm$ 0.002 2.717 $\pm$ 0.002 3.080 $\pm$ 0.016 0.881 $\pm$ 0.002
Scaffold [26] 0.995 $\pm$ 0.001 0.602 $\pm$ 0.001 0.703 $\pm$ 0.002 -8.116 $\pm$ 0.002 0.695 $\pm$ 0.001 2.728 $\pm$ 0.008 2.968 $\pm$ 0.038 0.776 $\pm$ 0.001
Scaffold Generic [26] 0.994 $\pm$ 0.001 0.617 $\pm$ 0.001 0.710 $\pm$ 0.002 -8.179 $\pm$ 0.012 0.701 $\pm$ 0.000 2.645 $\pm$ 0.008 3.090 $\pm$ 0.029 0.801 $\pm$ 0.000
DrugImprover [22] 0.884 $\pm$ 0.005 0.432 $\pm$ 0.002 0.493 $\pm$ 0.005 -6.726 $\pm$ 0.007 0.506 $\pm$ 0.002 1.306 $\pm$ 0.010 2.057 $\pm$ 0.011 0.531 $\pm$ 0.002
Molsearch [40] 1.000 $\pm$ 0.001 0.616 $\pm$ 0.001 0.726 $\pm$ 0.002 -8.855 $\pm$ 0.040 0.686 $\pm$ 0.001 3.105 $\pm$ 0.006 2.452 $\pm$ 0.008 0.969 $\pm$ 0.001
MIMOSA [10] 0.985 $\pm$ 0.008 0.622 $\pm$ 0.001 0.734 $\pm$ 0.002 -8.800 $\pm$ 0.015 0.677 $\pm$ 0.004 3.105 $\pm$ 0.008 2.711 $\pm$ 0.010 0.959 $\pm$ 0.001
DrugEx v3 [24] 1.000 $\pm$ 0.001 0.524 $\pm$ 0.001 0.613 $\pm$ 0.001 -8.089 $\pm$ 0.013 0.583 $\pm$ 0.002 3.095 $\pm$ 0.005 3.932 $\pm$ 0.008 0.495 $\pm$ 0.001
@@
(PDBID: MMP [26] 0.998 $\pm$ 0.001 0.636 $\pm$ 0.000 0.731 $\pm$ 0.001 -8.465 $\pm$ 0.021 0.709 $\pm$ 0.001 2.599 $\pm$ 0.004 3.013 $\pm$ 0.013 0.845 $\pm$ 0.001
4DWQ) Similarity ($\geq$ 0.5) [26] 0.999 $\pm$ 0.001 0.626 $\pm$ 0.000 0.723 $\pm$ 0.001 -8.511 $\pm$ 0.012 0.713 $\pm$ 0.002 2.543 $\pm$ 0.002 3.082 $\pm$ 0.031 0.760 $\pm$ 0.000
Similarity ([0.5, 0.7)]) [26] 0.999 $\pm$ 0.001 0.622 $\pm$ 0.001 0.718 $\pm$ $0$8.486 $\pm$ 0.021 0.713 $\pm$ 0.003 2.542 $\pm$ 0.005 3.101 $\pm$ 0.005 0.740 $\pm$ 0.001
Similarity ($\geq$ 0.7) [26] 0.999 $\pm$ 0.001 0.639 $\pm$ 0.000 0.734 $\pm$ 0.001 -8.496 $\pm$ 0.009 0.718 $\pm$ 0.001 2.628 $\pm$ 0.001 2.868 $\pm$ 0.003 0.875 $\pm$ 0.002
Scaffold [26] 0.998 $\pm$ 0.001 0.609 $\pm$ 0.001 0.718 $\pm$ $0$8.508 $\pm$ 0.026 0.711 $\pm$ 0.000 2.627 $\pm$ 0.002 2.803 $\pm$ 0.010 0.735 $\pm$ 0.002
Scaffold Generic [26] 0.998 $\pm$ 0.001 0.625 $\pm$ 0.001 0.722 $\pm$ $0$8.544 $\pm$ 0.009 0.722 $\pm$ 0.002 2.551 $\pm$ 0.010 2.898 $\pm$ 0.005 0.768 $\pm$ 0.004
DrugImprover [22] 0.920 $\pm$ 0.008 0.478 $\pm$ 0.001 0.618 $\pm$ 0.002 -8.701 $\pm$ 0.037 0.486 $\pm$ 0.002 1.181 $\pm$ 0.010 2.026 $\pm$ 0.013 0.427 $\pm$ 0.001
Molsearch [40] 1.000 $\pm$ 0.001 0.625 $\pm$ 0.001 0.742 $\pm$ 0.001 -8.747 $\pm$ 0.009 0.719 $\pm$ 0.001 3.012 $\pm$ 0.004 2.273 $\pm$ 0.005 0.950 $\pm$ 0.001
MIMOSA [10] 0.989 $\pm$ 0.001 0.631 $\pm$ 0.001 0.749 $\pm$ 0.001 -8.972 $\pm$ 0.011 0.706 $\pm$ 0.003 3.080 $\pm$ 0.007 2.561 $\pm$ 0.008 0.945 $\pm$ 0.001
DrugEx v3 [24] 1.000 $\pm$ 0.001 0.592 $\pm$ 0.001 0.668 $\pm$ 0.001 -8.762 $\pm$ 0.010 0.583 $\pm$ 0.002 2.488 $\pm$ 0.005 5.827 $\pm$ 0.010 0.393 $\pm$ 0.001
@@
Table 1: Main results. A comparison of eight baselines including Original, six baselines from REINVENT {MMP, Similarity ($\geq 0.5$), Similarity $\in[0.5,0.7)$, Similarity $\geq 0.7$, Scaffold, Scaffold Generic}, DrugImprover, Molsearch, MIMOSA and different versions of ScaffoldGPT on various objectives based on 3CLPro and RTCB datasets. The top two results are highlighted as 1st and 2nd. Results are reported for 5 experimental runs.
APO makes policy gradient based on the advantage preference [22], which is defined as
and perform APO policy gradient with follows:
Stage 3. Token-level Controllable decoding generation.
Ultimately, the current LLM-based decoder focuses mainly on maximizing likelihood, neglecting specific metrics of interest. This approach limits its effectiveness in optimizing objectives that diverge from those in its training set, particularly in generating desired molecules. In this study, we introduce controllable decoding after fine-tuning with APO. We present a new approach, Top-N, to direct the generation process towards enhancements in the optimization objective, as detailed below:
$Y^{\star}$$\sim\left\{[\mathbf{y}_{<i},y_{i},\cdots,y_{T}]\right\},$ (12)
[…]
 $\text{~{}}\mathcal{A}_{\mathbf{y}_{<i}}=\left\{y_{1},\ldots,y_{n}\right\},y_{i}\in\mathcal{V},|\mathcal{A}|\leq k,$
 and $\text{~{}and }R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{g},N,R}\right)|_{S,p,k}}\right)$
$\geq R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{h},N,R}\right)|_{S,p,k}}\right),\forall g<h,$
@@
5 EXPERIMENTS
5.1 Experimental configuration
The language model. We utilize GPT-2-like Transformers for causal language modeling, employing the standard 11M Drug-like Zinc dataset for training. Entries with empty scaffold SMILES are excluded, and we adopt a 90/10 split for training and validation, respectively. The training process is structured into three phases: pretraining, fine-tuning, and decoding optimization, as outlined in Algorithm ScaffoldGPT (See appendix for more details).
Original                                 Scaffold improved 1 improved 2 improved 3
Molecule [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE String Cc1ccc(O)c(Nc2nnc (-c3ccccc3)c(=O)[nH]2)c1 Cc1ccc(-c2nnc(Nc3ccc (C(C)C)cc3)[nH]c2=O)cc1 COc1ccc(-c2nnc(Nc3cccc (C)c3)[nH]c2=O)cc1 Cc1ccc(-c2nnc(Nc3cccc (C(C)C)c3)[nH]c2=O)cc1
Scaffold O=c1[nH]c(Nc2ccccc2)nnc1-c1ccccc1 same same same
Similarity ()($\uparrow$)( ↑ ) - 0.826 0.911 0.866
Docking ()($\downarrow$)( ↓ ) -10.031 -11.478 $\checkmark$ -11.474 $\checkmark$ -11.087 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.646 0.762 $\checkmark$ 0.774 $\checkmark$ 0.762 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 2.390 2.298 $\checkmark$ 2.257 $\checkmark$ 2.356 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 2.590 4.007 $\checkmark$ 2.893 $\checkmark$ 4.007 $\checkmark$
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.618 0.759 $\checkmark$ 0.753 $\checkmark$ 0.754 $\checkmark$
Table 2: One optimization example from cancer benchmark. Every generated molecules retains the scaffold, with all desired properties improved compared to the original. $\checkmark$ indicates improved property.
Baselines. In this work, we employ baseline models including DrugImprover [22], which utilizes an LSTM-based generator with APO fine-tuning, Molsearch [40], a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization, MIMOSA [10], a sampling-based method leveraging graph-based molecular optimization and DrugEx v3 [24], which leverages transformer-based reinforcement learning for scaffold-guided drug optimization. Additionally, we incorporate the model proposed by He et al. [15, 16], Loeffler et al. [26], which trains a transformer to adhere to the Matched Molecular Pair (MMP) guidelines [17, 43]. Specifically, given a set $\{\{X,Y,Z\}\}$, where $X$ represents the source molecule, $Y$ denotes the target molecule, and $Z$ signifies the property change between $X$ and $Y$, the model learns a mapping from $\{X,Z\}\in\mathcal{X}\times\mathcal{Z}\implies Y\in\mathcal{Y}$ during training. Here, $\mathcal{X}\times\mathcal{Z}$ denotes the input space, and $\mathcal{Y}$ denotes the target space. They defined six different types of property changes for $Z$, including MMP for user-specified alterations, various similarity thresholds, and scaffold-based modifications where molecules share the same scaffold or a generic scaffold. More specifically,
  • MMP: there exists user-specified desirable property changes between molecule $X$ and $Y$.
  • Similarity $\geq 0.5$: the tanimoto similarity between molecule $X$ and $Y$ is larger than 0.5.
  • Similarity $\geq 0.7$: the tanimoto similarity between molecule $X$ and $Y$ is larger than 0.7.
  • Scaffold: molecule $X$ and $Y$ share same scaffold.
  • Scaffold generic: molecule $X$ and $Y$ share same generic scaffold.
Dataset. We employ, from the most recent Cancer and COVID dataset of Liu et al. [22], 1 million compounds from the ZINC15 dataset docked to the 3CLPro (PDB ID: 7BQY) protein associated with SARS-CoV-2 and the RTCB (PDB ID: 4DWQ) human cancer protein.
Critics and evaluation metric. In this study, we evaluate the efficacy of ScaffoldGPT in generating molecules with desirable attributes within the context of pharmaceutical drug discovery. We leverage the RDKit [19] chemoinformatics package and employ various performance metrics as follows: Validity: It measures if the generated SMILES is valid in syntax. Druglikeness: The druglikeness metric measures the likelihood of a molecule being a suitable candidate for drug development. Solubility: This metric assesses the likelihood of a molecule’s ability to mix with water, commonly referred to as the water-octanol partition coefficient (LogP). Synthetizability: This parameter quantifies the ease (score of 1) or difficulty (score of 10) associated with synthesizing a given molecule [7]. Docking Score: The docking score assesses the drug’s potential to bind and inhibit the target site. To enable efficient computation, we employ a docking surrogate model (See Appendix A.4) to output this score. Similarity: We use Tanimoto similarity to evaluate the similarity between original SMILES and generated SMILES. Average Top 10% Norm Reward: It is the average of the normalized reward of the top 10% of molecules based on their average normalized reward. Average Norm Reward: It is the average of the normalized values of the docking score, druglikeness, synthesizability, solubility, and similarity across all valid molecules. This is the most important metric.
@@ 5.1 Experimental configuration
In contrast to the current state-of-the-art approach, REINVENT4, which pre-trains a Transformer with constraints on Tanimoto similarity, their method falls short in achieving drug improvement as it overlooks the optimization of multiple pharmaceutical properties. Therefore, Table 1 reveals that although REINVENT4 achieved high similarity, the generated molecules often failed to surpass the original ones. ScaffoldGPT, on the other hand, employs the APO reinforcement learning algorithm to fine-tune the pre-trained generative model and utilizes the Top-N decoding optimization strategy. These approaches ensure improvements aligned with multiple pharmaceutical objectives and enable ScaffoldGPT to successfully enhance the original drug across various pharmaceutical properties while maintaining a high Tanimoto similarity.
@@
1024 SMILES Details
1024 $>100$ length, $>50$ scaffold
Validity
One phase 0.49
Two phases 0.67
Table 3: Ablation study of one-phase vs two-phases
@@
Table 4: Ablation study of Top-N, Top-P, Top-K and Top-PK sampling strategies. The top result is highlighted as 1st. Top-N outperforms others in most of the metrics.
In this section, we present ablation studies that underscore the necessity and effectiveness of each component of ScaffoldGPT. These components complement each other, substantially enhancing overall performance.
@@
In two-phase incremental pretraining, the intuition behind the first phase lies in training on critical keywords as knowledge pieces, reinforcing the memory of these key terms, particularly in longer sequences. We conducted an ablation study comparing two-phase incremental training with one-phase training in Table 3. The results showed that the two-phase approach significantly improves validity compared to one-phase training, demonstrating the effectiveness of the incremental training method.

### v2

@@ 4 The ScaffoldGPT Algorithm
 $\text{~{}and }R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{g},N,R}\right)|_{S,p,k}}\right)\geq R\left({\text{{{BON}}}\left({\mathbf{y}_{<i}\circ y_{h},N,R}\right)|_{S,p,k}}\right),\forall g<h,$
Remark 4.1.
5 EXPERIMENTS
5.1 Experimental configuration
The language model. We use GPT-2-like Transformers for causal language modeling, employing the standard 11M Drug-like Zinc dataset for training. Entries with empty scaffold SMILES are excluded, and we adopt a 90/10 split for training and validation, respectively. The training process is structured into three phases: pretraining, fine-tuning, and decoding optimization, as outlined in Algorithm ScaffoldGPT (See appendix for more details).
Baselines. We compare against baseline models: DrugImprover (Liu et al., 2023a), which utilizes an LSTM-based generator with APO fine-tuning; Molsearch (Sun et al., 2022), a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization; MIMOSA (Fu et al., 2021), a sampling-based method leveraging graph-based molecular optimization; and DrugEx v3 (Liu et al., 2023b), which leverages transformer-based reinforcement learning for scaffold-guided drug optimization. Additionally, we incorporate the model proposed by He et al. (2021; 2022), Loeffler et al. (2024), which trains a transformer to adhere to the Matched Molecular Pair (MMP) guidelines (Kenny and Sadowski, 2005, Tyrchan and Evertsson, 2017). Specifically, given a set $\{\{X,Y,Z\}\}$, where $X$ represents the source molecule, $Y$ denotes the target molecule, and $Z$ signifies the property change between $X$ and $Y$, the model learns a mapping from $\{X,Z\}\in\mathcal{X}\times\mathcal{Z}\implies Y\in\mathcal{Y}$ during training. Here, $\mathcal{X}\times\mathcal{Z}$ denotes the input space, and $\mathcal{Y}$ denotes the target space. They defined six different types of property changes for $Z$, including MMP for user-specified alterations, various similarity thresholds, and scaffold-based modifications where molecules share the same scaffold or a generic scaffold. More specifically,
  • MMP: there exists user-specified property changes between molecule $X$ and $Y$.
  • Similarity $\geq 0.5$: tanimoto similarity between molecule $X$ and $Y$ is larger than 0.5.
  • Similarity $\geq 0.7$: tanimoto similarity between molecule $X$ and $Y$ is larger than 0.7.
  • Scaffold: molecule $X$ and $Y$ share same scaffold.
  • Scaffold generic: molecule $X$ and $Y$ share same generic scaffold.
@@ 5.1 Experimental configuration
(PDBID: MMP (Loeffler et al., 2024) 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ $0$8.259 $\pm$ 0.004 0.691 $\pm$ 0.001 2.682 $\pm$ 0.004 3.109 $\pm$ 0.020 0.862 $\pm$ 0.000
7BQY) Similarity ($\geq$ 0.5) (Loeffler et al., 2024) 0.995 $\pm$ 0.001 0.615 $\pm$ 0.000 0.706 $\pm$ 0.001 -8.165 $\pm$ 0.024 0.697 $\pm$ 0.004 2.621 $\pm$ 0.006 3.180 $\pm$ 0.029 0.782 $\pm$ 0.001
Similarity ([0.5, 0.7)]) (Loeffler et al., 2024) 0.995 $\pm$ 0.001 0.612 $\pm$ 0.001 0.701 $\pm$ 0.001 -8.187 $\pm$ 0.010 0.691 $\pm$ 0.001 2.611 $\pm$ 0.009 3.240 $\pm$ 0.014 0.756 $\pm$ 0.003
Similarity ($\geq$ 0.7) (Loeffler et al., 2024) 0.995 $\pm$ 0.001 0.628 $\pm$ 0.001 0.718 $\pm$ 0.001 -8.214 $\pm$ 0.002 0.691 $\pm$ 0.002 2.717 $\pm$ 0.002 3.080 $\pm$ 0.016 0.881 $\pm$ 0.002
Scaffold (Loeffler et al., 2024) 0.995 $\pm$ 0.001 0.602 $\pm$ 0.001 0.703 $\pm$ 0.002 -8.116 $\pm$ 0.002 0.695 $\pm$ 0.001 2.728 $\pm$ 0.008 2.968 $\pm$ 0.038 0.776 $\pm$ 0.001
Scaffold Generic (Loeffler et al., 2024) 0.994 $\pm$ 0.001 0.617 $\pm$ 0.001 0.710 $\pm$ 0.002 -8.179 $\pm$ 0.012 0.701 $\pm$ 0.000 2.645 $\pm$ 0.008 3.090 $\pm$ 0.029 0.801 $\pm$ 0.000
DrugImprover (Liu et al., 2023a) 0.884 $\pm$ 0.005 0.432 $\pm$ 0.002 0.493 $\pm$ 0.005 -6.726 $\pm$ 0.007 0.506 $\pm$ 0.002 1.306 $\pm$ 0.010 2.057 $\pm$ 0.011 0.531 $\pm$ 0.002
Molsearch (Sun et al., 2022) 1.000 $\pm$ 0.001 0.616 $\pm$ 0.001 0.726 $\pm$ 0.002 -8.855 $\pm$ 0.040 0.686 $\pm$ 0.001 3.105 $\pm$ 0.006 2.452 $\pm$ 0.008 0.969 $\pm$ 0.001
MIMOSA (Fu et al., 2021) 0.985 $\pm$ 0.008 0.622 $\pm$ 0.001 0.734 $\pm$ 0.002 -8.800 $\pm$ 0.015 0.677 $\pm$ 0.004 3.105 $\pm$ 0.008 2.711 $\pm$ 0.010 0.959 $\pm$ 0.001
DrugEx v3 (Liu et al., 2023b) 1.000 $\pm$ 0.001 0.524 $\pm$ 0.001 0.613 $\pm$ 0.001 -8.089 $\pm$ 0.013 0.583 $\pm$ 0.002 3.095 $\pm$ 0.005 3.932 $\pm$ 0.008 0.495 $\pm$ 0.001
@@
(PDBID: MMP (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.636 $\pm$ 0.000 0.731 $\pm$ 0.001 -8.465 $\pm$ 0.021 0.709 $\pm$ 0.001 2.599 $\pm$ 0.004 3.013 $\pm$ 0.013 0.845 $\pm$ 0.001
4DWQ) Similarity ($\geq$ 0.5) (Loeffler et al., 2024) 0.999 $\pm$ 0.001 0.626 $\pm$ 0.000 0.723 $\pm$ 0.001 -8.511 $\pm$ 0.012 0.713 $\pm$ 0.002 2.543 $\pm$ 0.002 3.082 $\pm$ 0.031 0.760 $\pm$ 0.000
Similarity ([0.5, 0.7)]) (Loeffler et al., 2024) 0.999 $\pm$ 0.001 0.622 $\pm$ 0.001 0.718 $\pm$ $0$8.486 $\pm$ 0.021 0.713 $\pm$ 0.003 2.542 $\pm$ 0.005 3.101 $\pm$ 0.005 0.740 $\pm$ 0.001
Similarity ($\geq$ 0.7) (Loeffler et al., 2024) 0.999 $\pm$ 0.001 0.639 $\pm$ 0.000 0.734 $\pm$ 0.001 -8.496 $\pm$ 0.009 0.718 $\pm$ 0.001 2.628 $\pm$ 0.001 2.868 $\pm$ 0.003 0.875 $\pm$ 0.002
Scaffold (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.609 $\pm$ 0.001 0.718 $\pm$ $0$8.508 $\pm$ 0.026 0.711 $\pm$ 0.000 2.627 $\pm$ 0.002 2.803 $\pm$ 0.010 0.735 $\pm$ 0.002
Scaffold Generic (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.625 $\pm$ 0.001 0.722 $\pm$ $0$8.544 $\pm$ 0.009 0.722 $\pm$ 0.002 2.551 $\pm$ 0.010 2.898 $\pm$ 0.005 0.768 $\pm$ 0.004
DrugImprover (Liu et al., 2023a) 0.920 $\pm$ 0.008 0.478 $\pm$ 0.001 0.618 $\pm$ 0.002 -8.701 $\pm$ 0.037 0.486 $\pm$ 0.002 1.181 $\pm$ 0.010 2.026 $\pm$ 0.013 0.427 $\pm$ 0.001
Molsearch (Sun et al., 2022) 1.000 $\pm$ 0.001 0.625 $\pm$ 0.001 0.742 $\pm$ 0.001 -8.747 $\pm$ 0.009 0.719 $\pm$ 0.001 3.012 $\pm$ 0.004 2.273 $\pm$ 0.005 0.950 $\pm$ 0.001
MIMOSA (Fu et al., 2021) 0.989 $\pm$ 0.001 0.631 $\pm$ 0.001 0.749 $\pm$ 0.001 -8.972 $\pm$ 0.011 0.706 $\pm$ 0.003 3.080 $\pm$ 0.007 2.561 $\pm$ 0.008 0.945 $\pm$ 0.001
DrugEx v3 (Liu et al., 2023b) 1.000 $\pm$ 0.001 0.592 $\pm$ 0.001 0.668 $\pm$ 0.001 -8.762 $\pm$ 0.010 0.583 $\pm$ 0.002 2.488 $\pm$ 0.005 5.827 $\pm$ 0.010 0.393 $\pm$ 0.001
@@
Table 1: Main results. A comparison of eight baselines including Original, six baselines from REINVENT {MMP, Similarity ($\geq 0.5$), Similarity $\in[0.5,0.7)$, Similarity $\geq 0.7$, Scaffold, Scaffold Generic}, DrugImprover, Molsearch, MIMOSA and different versions of ScaffoldGPT on various objectives based on 3CLPro and RTCB datasets. The top two results are highlighted as 1st and 2nd. represents the top-priority target objective. Results are reported for five experimental runs.
Dataset. We employ, from the most recent Cancer and COVID dataset of Liu et al. (2023a), 1 million compounds from the ZINC15 dataset docked to the 3CLPro (PDB ID: 7BQY) protein associated with SARS-CoV-2 and the RTCB (PDB ID: 4DWQ) human cancer protein.
Critics and evaluation metric. In this study, we evaluate the efficacy of ScaffoldGPT in generating molecules with desirable attributes within the context of pharmaceutical drug discovery. We leverage the RDKit (Landrum et al., ) chemoinformatics package and employ various performance metrics as follows: Validity measures if the generated SMILES is valid in syntax. Druglikeness measures the likelihood of a molecule being a suitable candidate for drug development. Solubility assesses the likelihood of a molecule’s ability to mix with water, commonly referred to as the water-octanol partition coefficient (LogP). Synthetizability quantifies the ease (score of 1) or difficulty (score of 10) associated with synthesizing a given molecule (Ertl and Schuffenhauer, 2009). Docking Score assesses the drug’s potential to bind and inhibit the target site. To enable efficient computation, we employ a docking surrogate model (See Appendix A.4) to output this score. Similarity: We use Tanimoto similarity to evaluate the similarity between original SMILES and generated SMILES. Average Top 10% Norm Reward is the average of the normalized reward of the top 10% of molecules based on their average normalized reward. Average Norm Reward is the average of the normalized values of the docking score, druglikeness, synthesizability, solubility, and similarity across all valid molecules. This is the most important metric. Evaluations are based on a sample of 1,280 molecules.
@@
A similarity score that is too high results in low structural diversity, while a score that is too low suggests the molecules have drifted too far from the original. Neither extreme is desirable. Our goal is to achieve a balanced level of similarity with meaningful variation. In this work, the primary optimization objective is the average normalized reward.
@@
In contrast to the current state-of-the-art approach, REINVENT4, which pre-trains a Transformer with constraints on Tanimoto similarity, their method falls short in achieving drug optimization as it overlooks the optimization of multiple pharmaceutical properties. Therefore, Table 1 reveals that although REINVENT4 achieved high similarity, the generated molecules often failed to surpass the original ones. ScaffoldGPT, on the other hand, employs the APO reinforcement learning algorithm to fine-tune the pre-trained generative model and utilizes the Top-N decoding optimization strategy. These approaches ensure improvements aligned with multiple pharmaceutical objectives and enable ScaffoldGPT to successfully enhance the original drug across various pharmaceutical properties while maintaining a high Tanimoto similarity.
@@
Table 2: Ablation study of Top-N, Top-P, Top-K and Top-PK sampling strategies. The top result is highlighted as 1st. represents the top-priority target objective. Evaluations are based on five random seeds. Top-N outperforms others in most of the metrics.
We next present ablation studies that underscore the necessity and effectiveness of each component of ScaffoldGPT. These components complement each other, substantially enhancing overall performance.
1024 SMILES Details
1024 $>100$ length, $>50$ scaffold
Validity
One phase 0.57
Two phases 0.68
Figure 2: Ablation study of one-phase vs two-phases. Evaluations are based on five random seeds.
@@
In two-phase incremental pretraining, the intuition behind the first phase lies in training on critical keywords as knowledge pieces, reinforcing the memory of these key terms, particularly in longer sequences. We conducted an ablation study comparing our novel two-phase incremental training with conventional one-phase training in Table 2. To ensure a fair comparison in terms of total training epochs, we trained for 10 epochs in the conventional single-phase setting, and for five epochs per phase in our two-phase setting. The results showed that the two-phase approach improves validity compared to one-phase training, demonstrating the effectiveness of the incremental training method.